e^x2-1=5 (Round to 4 decimals)
Answer by lwsshak3(11628)

e^x2-1=5 (Round to 4 decimals)
e^x2=6
Take log of both sides
x^2*lne=ln6
lne=1
x^2=ln6
x=±√ln6≈±1.3386